### v2

# Multimodal ArXiv: A Dataset for Improving Scientific Comprehension of Large Vision-Language Models

Abstract
Large vision-language models (LVLMs), exemplified by GPT-4V, excel across diverse tasks involving concrete images from natural scenes. However, their ability to interpret abstract figures, such as geometry shapes and scientific plots, remains limited due to a scarcity of training datasets in scientific domains. To fill this gap, we introduce Multimodal ArXiv, consisting of ArXivCap and ArXivQA, for enhancing LVLMs scientific comprehension. ArXivCap is a figure-caption dataset comprising 6.4M images and 3.9M captions sourced from 572K ArXiv papers spanning various scientific domains. Drawing from ArXivCap, we introduce ArXivQA, a question-answering dataset generated by prompting GPT-4V based on scientific figures. ArXivQA greatly enhances LVLMs’ mathematical reasoning capabilities, achieving a 10.4% absolute accuracy gain on a multimodal mathematical reasoning benchmark. Furthermore, employing ArXivCap, we devise four vision-to-text tasks for benchmarking LVLMs. Evaluation results with state-of-the-art LVLMs underscore their struggle with the nuanced semantics of academic figures, with domain-specific training yielding substantial performance gains. Our error analysis uncovers misinterpretations of visual context, recognition errors, and the production of overly simplified captions by current LVLMs, shedding light on future improvements. [^1] Project page: https://mm-arxiv.github.io.
Multimodal ArXiv: A Dataset for Improving Scientific Comprehension of Large Vision-Language Models
Lei Lithanks: Equal ${}^{\dagger}$, Yuqi $*{}^{*}$${}^{\dagger}$, Runxin ${}^{\ddagger}$, Peiyi ${}^{\ddagger}$ Xiachong ${}^{\dagger}$, Lingpeng ${}^{\dagger}$, Qi ${}^{\dagger}$$^{\dagger}$he University of Hong Kong ${}^{\ddagger}$eking University {nlp.lilei, runxinxu, wangpeiyi9979, xiachongfeng1996}@gmail.com wangyuqi@connect.hku.hk{lpk, liuqi}@cs.hku.hk
1 Introduction
Large vision-language models (LVLMs), which integrate large language models (LLMs) (Brown et al., 2020a; Touvron et al., 2023) with pre-trained vision encoders through cross-modal alignment training (Madureira, 2021; Liu et al., 2023b; Li et al., 2023d), have demonstrated remarkable perceptual and cognitive capabilities in processing concrete images from everyday scenes (OpenAI, 2023; Fu et al., 2023; Yang et al., 2023a). However, recent studies have exposed their limited comprehension of abstract figures, such as geometric shapes in multimodal mathematical reasoning (Lu et al., 2023) and scientific plots (Yue et al., 2023). The inadequacy of training datasets in scientific domains that involve complex reasoning with abstract figures is the main underlying cause.
To address this, we construct Multimodal ArXiv by utilizing the rich resources in preprints hosted on arXiv to improve the ability to understand scientific literature in LVLMs. We first curate ArXivCap, a diverse scientific figure-caption dataset. In contrast to previous scientific figure datasets, which consist of synthesized figures (Chen et al., 2020) or are restricted to simple captioning scenarios in the computer science domain (Hsu et al., 2021), our dataset is composed of figures extracted from academic papers across a range of domains. ArXivCap has 6.4M images and 3.9M captions from 572K papers. We also keep the subfigure structure, and titles of original papers, thereby supporting diverse evaluation tasks. We further instruct GPT-4V to generate 100K multiple-choice question-answering (QA) pairs for the figures in ArXivCap. The resulting ArXivQA dataset could naturally serve as a pivotal resource for improving the scientific reasoning abilities of LVLMs.
We validate the effectiveness of our Multimodal ArXiv from two dimensions: reasoning ability measured by QA accuracy and generation performance through captioning-like tasks. Our experiments demonstrate that ArXivQA brings a significant 10.4% absolute accuracy boost on the MathVista (Lu et al., 2023), a challenging benchmark for multimodal mathematical reasoning. Additionally, detailed analysis uncovers the relationship between paper domains and fine-grained task performance. Moreover, using ArXivCap, we define four generation tasks of varying complexity to benchmark the ability of LVLMs to comprehend scientific plots: (1) captioning a single academic figure, (2) generating overall summaries for multiple sub-figures, (3) in-context image captioning given previous figure-caption pairs, and (4) generating paper titles from figure-caption pairs. We examine various LVLMs, including open-source models as well as proprietary models including GPT-4V (OpenAI, 2023) and Bard (Google, 2023). Evaluation results reveal that despite that current LVLMs still face challenges generating faithful captions for scientific figures, in-domain training on our dataset yields substantial performance improvements across all four tasks. Manual error analysis underscores that LVLMs still suffer from misinterpretation of the visual context, recognition errors, and overly simplified captions, paving the way for future studies.
2 Related Work
Recent advancements in LVLMs have seen notable progress in model architecture, training paradigms, and dataset creation.
Model Architecture LVLMs typically comprise three core modules: (i) a vision encoder for image feature extraction, (ii) a modality alignment module to integrate visual features into the language model embedding space, and (iii) an LLM backbone for decoding multimodal context. CLIP (Radford et al., 2021) is widely used for image encoding, while LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023) serve as popular choices for LLMs. The alignment module varies from simple linear projections (Liu et al., 2023b; Zhu et al., 2023) to more complex architectures like gated cross-attention layers  substantiated by Flamingo and IDEFICS(Alayrac et al., 2022; Awadalla et al., 2023). Innovations such as the Q-Former module in BLIP2 (Li et al., 2023b) and instruction integration in InstructBLIP (Dai et al., 2023) further enhance alignment capabilities. Additionally, Fuyu-8B (Bavishi et al., 2023) introduces a novel framework mapping raw image pixels directly to the LLM embedding space.
Training Paradigms PaLI-X (Chen et al., 2023b) investigates scaling effects of both vision encoders and language models, highlighting the advantages of scaling both components. Qwen-VL (Bai et al., 2023b) increases input image resolution and explores different module unfreezing strategies. Alignment methodologies such as RLHF training (Ouyang et al., 2022), e.g., LLaVA-RLHF (Sun et al., 2023), and preference optimization through AI feedback (Li et al., 2023c) demonstrate improved alignment with human preferences.
[…]
3 Multimodal ArXiv
This section presents a detailed construction process of our Multimodal ArXiv dataset, consisting of two sets: ArXivCap (§3.1) and ArXivQA (§3.2).
Figure 1: Overview of our dataset curation process. Starting from the ArXiv paper source files, we ensure the paper quality by selecting papers according to publication records. Figure and caption pairs are extracted and then cleaned according to manually designed rules. ArXivQA is generated by prompting GPT-4V with a curated template.
3.1 ArXivCap
We outline the creation process of ArXivCap below and Figure 1 gives an overview.
Paper Filtering with Publication Type ArXivCap is extracted from ArXiv Clement et al. (2019), which is under CC-0 licence for modification and distribution. The raw files of papers posted on ArXiv tar files before June 2023 are downloaded. To ensure the quality of our dataset, we employ a rigorous selection process to filter potentially low-quality papers that might influence the figure-caption pair quality. Firstly, we retrieve meta-information for papers from Semantic Scholar Kinney et al. (2023), which contains the publication record for each paper. Papers with publication types JournalArticle, Conference, or Review are kept as we assume the peer-review process could ensure the overall figure-caption quality is satisfactory. We further treat papers with titles longer than 100 words, or with an abstract longer than 300 words according to common submission requirements and remove them.
Figure-Caption Pair Extraction Images and captions are extracted from the original LaTeX files by matching the syntax. We further use a robust tool ImageMagisk ImageMagick Studio LLC to convert images into JPEG format for easy processing. The extracted images and captions are stored in a designed chunk structure, which consists of either a single figure-caption pair or multiple figures with their respective sub-captions and a main caption for the overall description. This format is more consistent with the layout of academic papers, and Figure 2 illustrates the chunk structure.
Figure 2: Illustration of two types of figure-caption pairs. (Left) Single-Figure pair. (Right) Multiple-Figure caption pair has multiple sub-figures with corresponding sub-captions and an overall main caption.
Caption Cleaning and Image Filtering After a manual inspection of the initially collected dataset, we design several transformations to clean the captions and filter the images. Caption Cleaning: (i) Chunks with captions shorter than 5 words are removed; (ii) For captions with LaTex expressions such as math formulas and references, we apply the pylatexenc to transform the LaTeX to text with math formulas retained, citations to a special symbol <cit.>, references to <ref>. An illustration of caption cleaning can be found in Appendix A.1. Image Filtering: We remove images that are deemed to be problematic according to the following rules: (i) Images with an aspect ratio larger than 100; (ii) Images with the shortest edge shorter than 224 pixels; and (iii) Images with pixel numbers larger than the decompression bombs threshold.
After these processes, 100 pairs are sampled to perform an additional manual inspection, where we found all of these pairs contained clear images and correct caption descriptions. We provide visualized samples in Appendix A.3.
3.1.1 Statistics of ArXivCap
Field Number Average Len. Quartile of Len.
Title 572K 10.4 (8, 10, 12)
Abstract 572K 167.6 (126, 165, 207)
Main Caption 3.9M 47.6 (15, 35, 65)
Subcaption 1.0M 4.8 (2, 3, 5)
Chunk Caption 3.9M 48.8 (16, 36, 67)
Images 6.4M N / A N / A
Table 1: Word count statistics for title, abstract, captions, and Image number. Chunk caption refers to the combination of subcaptions and the main caption for a multiple-figure case.
Dataset Image Number Paper Number Image Category Domain Real Data
FigCAP (Chen et al., 2020) 219K N / A Bar, Line and Pie Charts N / A
SciCap (Yang et al., 2023b) 2.1M 295K Open-Category Computer Science and Machine Learning
M-Paper (Hu et al., 2023) 350K 48K Open-Category Mainly "Deep Learning"
ArXivCap (Ours) 6.4M 572K Open-Category Open-Domain
FigureQA (Kahou et al., 2017) 140K N / A Bar, Line and Pie Charts N / A
DVQA (Kafle et al., 2018) 300K N / A Bar Charts N / A
ArXivQA (Ours) 32K 16.6K Open-Category Open-Domain
Table 2: Comparison with previous scientific figure datasets. Our ArXivCap is the largest captioning dataset and our ArXivQA is the only QA dataset that covers a wide range of domains from real papers.
Table 1 lists the dataset statistics. ArXivCap consists of 572K papers, containing 6.4M high-quality images in total with 193M words. A word cloud illustration of captions can be found in the Appendix A.2. Figure 3 demonstrates the paper domain distribution extracted from ArXiv, where we find that our ArXivCap covers 32 paper domains, such as computer science, mathematics, physics, and economics. As shown in Table 2, compared with previous scientific figure datasets, our ArXivCap is the largest figure-caption dataset collected from real papers and covers a wide range of scientific domains, serving as a valuable resource for improving and benchmarking LVLMs.
Figure 3: Paper domain distribution of ArXivCap. See Table 8 in Appendix A for the full name of each domain.
3.2 ArXivQA
As our ArXivCap contains diverse images from scientific domains, we assume that learning to answer questions about these figures could boost scientific reasoning ability. Following the successful practice of LLaVA (Liu et al., 2023b), we adopt GPT-4V to generate instruction-tuning datasets for generating the QA pairs based on the figures extracted from scientific papers. Specifically, we design a prompting template to query GPT-4V for generating QA pairs based on 35K images randomly sampled from our ArXivCap. Table 9 in Appendix A.4 provides the template we used for the prompt. The generated pairs are parsed according to the format requirement and we discard the samples without options and rationales. There are 100K QA pairs after filtering the invalid samples. The dataset comprises questions with an average word count of 16.98 for the question text, and options for the question with an average word count of 31.86. On average, there are 4.20 options per question. Appendix A.3 provides samples from the ArXivQA dataset.
4 Experiments
In this section, we conduct experiments to (i) validate the effectiveness of ArXivQA (§4.1) and (ii) benchmark the LVLMs capability to comprehend scientific figures with our ArXivCap (§4.2).
4.1 Boosting LVLMs with ArXivQA
4.1.1 Experimental Settings
We adopt Qwen-VL-Chat (Bai et al., 2023b) as the backbone, and fine-tune it on our ArXivCap and ArXivQA datasets for three epochs with a learning rate of 1e-5 following the original paper. Models are evaluated on MathVista (Lu et al., 2023), a benchmark that requires fine-grained, deep visual understanding and compositional reasoning. MathVista contains 6,141 examples, consisting of five multimodal tasks Figure QA, Geometry Problem Solving, Math word problem, Text Book QA, and Visual QA. We select the multiple-choice questions in the benchmark to avoid the inconsistency of answer parsing. We compute the accuracy scores on the testmini set and adopt the provided prediction files for calculating the baseline performance.
4.1.2 Results
Model Figure QA Geometry Problem Solving Math Word Problem Textbook QA Visual QA ALL
${}^{\dagger}$ 41.4 22.0 18.2 34.6 44.6 33.7
${}^{\dagger}$ 41.4 19.9 45.5 45.8 57.6 39.3
${}^{\dagger}$ 44.0 26.7 40.9 45.8 44.6 39.3
Qwen-VL-Chat 48.3 19.1 22.7 46.7 57.6 40.0
 w/ ArXivCap 39.7 19.8 27.2 39.7 52.1 36.2
 w/ ArXivQA 44.8 34.0 27.3 70.0 64.1 50.2
 w/ ArXivQA + ArXivCap 44.0 37.6 27.3 68.2 63.0 50.4
${}^{\dagger}$ 38.8 51.1 27.3 64.5 51.1 50.0
${}^{\dagger}$ 52.6 51.8 54.5 83.2 66.3 61.9
Table 3: Evaluation on MathVista dataset. ArXivCap and ArXivQA together enhance Qwen-VL-Chat’s overall performance, surpassing that of the commercial model Bard. ${}^{\dagger}$ denotes results based on the original predictions from Lu et al. (2023). The best results are highlighted in bold, while the second-best scores are marked with underline.
As shown in Table 3, fine-tuning on our Multimodal ArXiv, especially on the ArXivQA dataset, consistently boosts the performance, helping the open-sourced Qwen-VL-Chat achieve a comparable overall MathVista reasoning performance. Due to the wide coverage of the scientific figures, the performance gain mainly comes from significantly improved Geometry Problem Solving, Textbook QA, and Visual QA tasks. For example, after fine-tuning on the ArXivQA dataset, the accuracy is increased from 46.7% to 70.0% and from 19.1% to 34.0% on Geometry Problem Solving and Textbook QA tasks, respectively. The improvement on Math Word Problem is marginal, where we think the domain-specific data augmentation can be further explored with a curated filtering dataset on our dataset (Gao et al., 2023). On the contrary, the accuracy of Figure QA deteriorates slightly compared with the original backbone model, which we attribute to the fact that most of the plots in the Figure QA evaluation are sampled from synthesized datasets such as DVQA (Kafle et al., 2018), exhibiting great gaps between real-world paper figures.
4.1.3 Analysis
We investigate how different subject domains affect mathematical reasoning ability using pairs of questions and answers (QA). We focus on six domains with more than 5K samples each. From each domain, we randomly choose a subset of 5K samples to ensure fairness in comparison. We then fine-tune the Qwen-VL-Chat base model using QA pairs from each domain and observe how it affects the model’s accuracy compared to its original state. Figure 4 demonstrates the relative accuracy changes (i.e., Accuracy after Fine-tuningOriginal $\frac{\text{Accuracy after Fine-tuning}}{\text{Original Accuracy}}-1$) after training the model on QA pairs from each domain. Our findings reveal several key points: (i) QA pairs from the Computer Science (CS) domain are highly effective for improving mathematical reasoning ability, achieving a notable 27.09% relative improvement. We attribute this to the compositional nature the the CS area. (ii) The most beneficial domain varies depending on the specific task. For instance, QA pairs from astrophysics domains enhance geometry problem-solving, while those from Condensed Matter improve performance in math word problems. (iii) Most domains hurt the Figure QA task. This suggests that synthetic Figure QA might not be the best benchmark for assessing realistic reasoning ability. These findings underscore the efficacy of generated QA pairs and offer valuable insights for future research, such as adjusting task-specific weights in the dataset accordingly.
Figure 4: Relative accuracy changes brought by the training on different domain ArXivQA samples.
4.2 Benchmarking LVLMs on ArXivCap
4.2.1 Evaluated Tasks
We design four vision-to-text tasks to benchmark LVLMs’ ability to comprehend scientific figures. Single-Figure Captioning Similar to the traditional image captioning setup (Lin et al., 2014; Hsu et al., 2021), single-figure captioning requires the model to generate a caption for the given single figure. The captions generated by the model are expected to encapsulate the nuanced details within these figures, including numbers and mathematical formulas, presenting a unique challenge for models to identify and articulate these elements accurately. Formally, given an image-caption pair $(I,C)$, the LVLM $\mathcal{M}$ is asked to generate the caption given an instruction prompt $P_{s}$ to hint the goal of scientific captioning:
$\hat{C}=\mathcal{M}(I,P_{s}),$ (1)
where $\hat{C}$ would be evaluated according to the ground-truth $C$.
Multiple-Figure Captioning We introduce a more intricate challenge involving applying reasoning across multiple images. This task, termed Multiple-Figure Captioning, necessitates the model to craft a comprehensive summary caption for subfigures. As exemplified in Figure 2, the model is tasked with generating an overarching caption for two or more subfigures, leveraging visual clues to draw comparisons and formulate summary captions. Formally, given a list of figures $L=\left(I_{1},\ldots,I_{n}\right)$, the model is to generate the ground-truth main caption $C$ by considering all the semantics in the figures with a task prompt $P_{m}$:
$\hat{C}=\mathcal{M}(L,P_{m})=\mathcal{M}(I_{1},\ldots,I_{n},P_{m}).$ (2)
Contextualized Image Captioning Inspired by the evolving in-context learning capabilities of LLMs (Brown et al., 2020b; Dong et al., 2022), we introduce a contextualized captioning task to examine the in-context learning ability of LVLMs. In this task, the model is presented with a set of figure-caption pairs, and its goal is to generate a caption for a given image based on the provided demonstrations. Given a sequential image-captions pairs $S=\{(I_{i},C_{i})\}_{i=1}^{n}$ consisting of $n$ pairs of image $I_{i}$ and the corresponding $C_{i}$, the contextualized image captioning task can be formalized as follows:
$\hat{C_{n}}=\mathcal{M}(I_{1},C_{1},\ldots,I_{n-1},C_{n-1},I_{n},P_{c}).$ (3)
The model is supposed to leverage the context history to enhance the accuracy and coherence of the generated caption.
Title Generation This task requires a nuanced understanding of figures and captions to distill essential observations into a high-level summary of the presented results for LVLMs. Specifically, instead of producing the captions for the figures, this task requires the model to connect different figures and corresponding captions to infer the paper title. Let $S=\{(I_{i},C_{i})\}_{i=1}^{m}$ be a sequence of $m$ figure-caption pairs in the extracted paper. Note that $I_{i}$ could be a single figure or a multiple-figure, and we reuse $I_{i}$ for simplicity here. The title generation asks $\mathcal{M}$ to generate the title for the paper given a task prompt $P_{t}$:
$\hat{T}=\mathcal{M}(I_{1},C_{1},\ldots,I_{m},C_{m},P_{t}).$ (4)
The prediction $\hat{T}$ is evaluated by comparing it to the original title $T$.
4.2.2 Experimental Settings
Model BLEU-2 ROUGE-L BERT-S
BLIP-2-OPT-6.7B 2.1 7.1 81.1
InstructBLIP-Vicuna7B 3.7 10.1 83.3
LLaVA-1.5-7B 2.3 10.6 83.0
LLaVA-1.5-13B 2.6 10.7 83.3
OpenFlamingo-9B 5.7 9.9 82.4
IDEFICS-Instruct-9B 2.5 9.1 83.5
Qwen-VL-Chat-7B 4.4 11.1 81.8
 w/ ArXivCap 8.9 15.8 83.3
Bard 3.4 12.2 82.2
GPT-4V 5.7 14.6 83.2
Table 4: Evaluation results of single figure captioning. Grey results are obtained from a 200-sample subset. Despite most LVLMs struggling to produce high-quality captions of scientific figures, training with ArXivCap significantly boosts the performance.
Evaluated Models We select various LVLMs covering different architectures. (1) LVLMs designed for dealing with a single image, BLIP2-OPT-6.7B (Li et al., 2023b) , InstructBLIP-Vicuna7B (Dai et al., 2023), LLaVA-1.5-7B/13B (Liu et al., 2023a). Due to the ability limitation, we only benchmark these models on the single image captioning task; (2) LVLMs capable of handling interleaved text-image inputs, such as OpenFlamingo-9B (Alayrac et al., 2022; Awadalla et al., 2023), IDEFICS-Instruct-9B (Laurençon et al., 2023), Qwen-VL-Chat-7B (Bai et al., 2023b). These models are evaluated on all the tasks we proposed; (3) Proprietary models such as Bard and GPT-4V. Due to the API limit of these models, we evaluate these models on a randomly sampled subset consisting of 200 instances for evaluation, with corresponding scores colored in grey. Details of evaluated models and the task prompts used are provided in Appendix B.
Training Settings To investigate whether in-domain training can enhance the model’s ability, we train the Qwen-VL-Chat-7B for three epochs and adopt the AdamW optimizer (Kingma and Ba, 2015) with a learning rate of 1e-5 and weight decay set to 0.05, following the original setup of Qwen-VL-Chat. We set the maximum figure number in a sample to $4$ to fit the input length limit. The training costs 70 hours with 8 NVIDIA A100 GPUs.
Metrics We adopt BLEU-2 (Papineni et al., 2002), ROUGE-L (Lin, 2004) and BERT-Score (Zhang et al., 2020) as the automatic evaluation metrics. A manual error analysis is performed to supplement the automatic evaluation (§5).
Model Multiple-Figure Captioning Contextualized Captioning Title Generation
BLEU-2 ROUGE-L BERT-S BLEU-2 ROUGE-L BERT-S BLEU-2 ROUGE-L BERT-S
OpenFlamingo-9B 3.7 11.3 81.9 20.0 20.5 83.7 2.7 17.7 82.7
IDEFICS-Instruct-9B 3.6 10.8 82.8 20.7 22.6 85.7 3.5 18.4 85.8
Qwen-VL-Chat-7B 3.0 7.2 79.7 17.0 22.1 85.0 2.6 15.8 85.1
 w/ ArXivCap 10.6 18.0 83.6 16.1 21.2 84.8 6.7 23.5 86.8
GPT-4V 5.7 14.7 83.0 9.6 21.1 84.8 3.4 19.3 86.0
Table 5: Evaluation results of three newly defined tasks. The best results are highlighted in bold.
4.2.3 Results
Results of Single-Figure Captioning The evaluation results for the single figure captioning task are presented in Table 4. Despite achieving almost perfect performance on conventional image captioning tasks such as MSCOCO (Lin et al., 2014), LVLMs encounter challenges when applied to academic figures. Among LVLMs, Qwen-VL-Chat emerges as the top-performing open-source candidate, surpassing its counterparts. Notably, GPT-4V excels in generating captions compared to the competing model, Bard. Moreover, continuous training on our dataset yields a significant performance boost for this task. For instance, fine-tuning results in a notable increase in the BLEU-2 score from 4.4 to 8.9, indicating a promising avenue for enhancing academic figure comprehension through domain-specific training.
Results of Multiple-Figure Captioning As shown in the first block of table 5, similar to single-figure captioning, multiple-image captioning poses a challenge for current open-source LVLMs. For instance, Qwen-VL-Chat achieves only a 3.0 BLEU-2 and a 7.2 ROUGE-L score on this task, considerably lower than its performance in single-figure captioning. In contrast, GPT-4V consistently demonstrates proficiency in both tasks, suggesting a balanced ability to capture semantics across multiple images. Notably, training on our ArXivCap dataset yields more pronounced improvements for this task, culminating in Qwen-VL-Chat even surpassing the performance of the GPT-4V model. This enhancement underscores the pivotal role of our dataset in facilitating LVLMs to enhance reasoning capabilities over multiple images, leading to more effective summarization of scientific figures.
Results of Contextualized Captioning In the middle block of Table 5, we find that IDEFICS-Instruct-9B achieves the best performance on this task. This achievement is largely attributed to its remarkable proficiency in leveraging contextual cues, stemming from its extensive pre-training involving interleaved image-text pairs (Laurençon et al., 2023). Interestingly, fine-tuning on ArXivCap results in marginal performance declines across all metrics, with GPT-4V achieving the lowest scores as well. This phenomenon can be attributed to the tendency of sequential captions to exhibit similar patterns, thereby favoring models that effectively leverage contextual cues. We further perform a more challenging evaluation by randomly shuffling the order of figure-caption pairs in the context. As shown in Table 6, the fine-tuned model demonstrates more robust captioning results under this setting, evidenced by the slight 8% drop on ROUGE-L compared to the 31% of the original model.
Model BLEU-2 ($\Delta\downarrow$) ROUGE-L ($\Delta\downarrow$)
Qwen-VL-Chat-7B 17.0 22.1
 + shuffle order 12.0 (29.4%) 15.1 (31.7%)
Qwen-VL-Chat-7B w/ ArXivCap 16.1 21.2
 + shuffle order 14.1 (12.4%) 19.5 (8.0%)
Table 6: Contextualized captioning performance is influenced by the order. After tuning on the ArXivCap, the model is more robust to the order of the history captions.
Results of Title Generation The results are presented in the last block of Table 5. Notably, the title generation task poses a formidable challenge, evident in the significantly lower overall BLEU-2 score compared to the captioning tasks. This suggests the inherent difficulty in generating precise predictions for paper titles. A contrasting picture emerges when considering the ROUGE-L and BERT-Score metrics, which either closely align or surpass the performance on captioning tasks. This underscores the model’s proficiency in producing semantic-related results given the presented figures. Consistent with the previous two tasks, fine-tuning the model on our dataset yields substantial enhancements for the title generation task. The BLEU-2 score jumps impressively from 2.6 to 6.7, while the ROUGE-L score sees a commendable increase from 15.8 to 23.5. These findings highlight the challenge of title generation for current LVLMs and the effectiveness of our dataset in improving the model’s capability to generate accurate titles.
4.2.4 Analysis
Figure 5: Manual analysis of the generated captions.
Manual Evaluation We conduct a manual inspection for single-figure captioning results. To ensure a more informed evaluation, we focus on a paper from the CS domain, leveraging our domain knowledge to assess caption quality better. The quality of generated captions is assessed by scrutinizing the figure, the ground-truth caption, the paper title, and the abstract. We categorize captions into the following quality types according to our preliminary inspection: (1) Acceptable, where captions accurately encapsulate the scientific figure’s essence, aligning with the intended information of the ground-truth; (2) Over Simplification, instances where the model oversimplifies content, offering a broad overview while neglecting specific details and nuances present in the ground truth; (3) Recognition Error, where the model inaccurately recognizes and describes key visual and textual elements in the scientific figure, such as colors, numerical values, or textual context; and (4) Contextual Misinterpretation, where the model misinterprets the specific context of the scientific figure, resulting in captions relevant in a generic sense but inaccurate for the given figure. Visualized generated captions of different types are shown in Figure 14 of Appendix C.1. The results of 100 manually examined captions are depicted in Figure 5, revealing that only 16% of captions are deemed acceptable when compared to human-written ones. Among unsatisfactory captions, contextual misinterpretation emerges as the dominant issue, suggesting a need for incorporating more contextual information. Oversimplification is another concern, with generic captions identified. Additionally, 23% and 19% of examined captions suffer from the oversimplification issue and recognition errors in reported numbers/texts in the caption, respectively. The former is attributed to the highly frequent simple caption in the training dataset and the latter issue could be addressed through potential integration with OCR results. Our manual evaluation suggests future efforts may benefit from incorporating additional context clues, such as paper metadata, improving the model’s fundamental perception abilities, and utilizing external information.
Figure 6: ArXivQA enables the model not only to answer questions related to scientific figures in papers (left) but also to improve mathematical understanding ability (right). The model not only selects correct options but also gives reasonable rationale.
Case Study We conducting case studies to illuminate the tuning effects facilitated by our ArXivQA dataset. In the left segment of Figure 6, ArXivQA helps the model accurately answer a question related to the presented bar plot. The right part in Figure 6 demonstrates that ArXivQA can enhance algebraic reasoning abilities. Here, a question involving the derivative of a function is correctly answered, accompanied by a lucid reasoning rationale. Figure 15 in Appendix C.2 highlights a challenging geometry problem where both models generate hallucinated outputs. These illustrative cases collectively affirm the efficacy of our dataset.
5 Conclusion
Our work introduces Multimodal ArXiv, comprising ArXivCap and ArXivQA, aims at advancing the scientific comprehension of LVLMs. Experiments show that fine-tuning on ArXivQA notably enhances LVLMs’ mathematical reasoning capabilities. Moreover, our comprehensive evaluations across four vision-to-text tasks on ArXivCap underscore the challenges in understanding scientific figures for LVLMs, while highlighting the substantial improvements achieved by in-domain training. Our error analysis offers valuable insights for the ongoing development of LVLMs.
Limitations
Our study has several limitations worth noting. Firstly, our exploration may not encompass the full spectrum of LVLMs due to the rapid evolution of architectures and training methodologies. Nevertheless, we believe our dataset could still be effective for other LVLMs and the findings are generalizable. Secondly, our Multimodal ArXiv primarily sources from ArXiv papers due to its accessibility and open-sourced license, potentially overlooking the diversity of disciplines and data modalities present in scientific literature. Future research could incorporate a broader range of datasets and domains to enrich the coverage of scientific knowledge.
Appendix A Details of Multimodal ArXiv
A.1 Caption Cleaning
We apply a Python tool to clean the original caption and Table 7 illustrates the caption before and after cleaning.
Before Cleaning
After Cleaning
[…]
As Fig. <ref >except at $z\sim 6$ ($z=4.37$ in the EdS model).
Table 7: Caption before and after cleaning using pylatexenc.
A.2 Caption Word Cloud
We visualize the word cloud of captions in our ArXivCap dataset in Figure 7, it can be seen that the captions have a diverse vocabulary for describing the different figures in the academic papers.
Figure 7: Word cloud visualization of captions in our ArXivCap dataset.
A.3 Illustration Cases from ArXivCap and ArXivQA
We provide illustrated cases from our dataset for a better understanding. Figure 8 demonstrates a typical single figure-caption pair, and Figure 9 shows the multiple figure-caption case. Figure 10, 11, 12 and 13 illustrate the cases from our ArXivQA dataset, covering different figure types and containing diverse questions.
Figure 8: A single-figure caption pair in our ArXivCap dataset.
Figure 9: A multiple-figure caption pair in our ArXivCap dataset.
Figure 10: A case from ArXivQA.
Figure 11: A case from ArXivQA.
Figure 12: A case from ArXivQA.
Figure 13: A case from ArXivQA.
Domain Full Name
dg-ga Differential Geometry
acc-phys Accelerator Physics
solv-int Exactly Solvable and Integrable Systems
q-alg Quantum Algebra and Topology
atom-ph Atomic, Molecular and Optical Physics
alg-geom Algebraic Geometry
comp-gas Cellular Automata and Lattice Gases
supr-con Superconductivity
chem-ph Chemical Physics
mtrl-th Materials Theory
adap-org Adaptation, Noise, and Self-Organizing Systems
patt-sol Pattern Formation and Solitons
chao-dyn Chaotic Dynamics
cmp-lg Computation and Language
econ Economics
hep-lat High Energy Physics - Lattice
nucl-ex Nuclear Experiment
q-fin Quantitative Finance
math-ph Mathematical Physics
nucl-th Nuclear Theory
gr-qc General Relativity and Quantum Cosmology
hep-ex High Energy Physics - Experiment
hep-th High Energy Physics - Theory
nlin Nonlinear Sciences
hep-ph High Energy Physics - Phenomenology
q-bio Quantitative Biology
quant-ph Quantum Physics
eess Electrical Engineering and Systems Science
stat Statistics
astro-ph Astrophysics
physics Physics
cond-mat Condensed Matter
math Mathematics
cs Computer Science
Table 8: Name of each domain.
Multiple-choice Question Answer Pairs Generation for Scientific Figures Guideline The goal of this task is to create answerable multiple-choice questions based on figures from scientific papers, to improve the ability of a large vision language model. The questions should be challenging, and require college-level reasoning. The type of questions should be diverse. The question should be answerable based on the figure. The answer should be one of the answer choices. The answer choices should be plausible and challenging. Format Below is an example of the format of the input and output for the task. Input Figures: [Figures input in the task] Output Question: [Question] Answer Options: [Answer choices, a bullet list.] Correct Choice: [Correct answer choice, e.g., A] Rationale: [Rationale for the correct answer, explain why the answer is correct]
Table 9: Prompt used for GPT-4V to generate QA pairs based on scientific figures.
A.4 ArXivQA Prompting Template
The prompt used to query GPT-4V is provided in Table 9.
Appendix B Evaluation Details
B.1 Details of Evaluated Models
BLIP2 (Li et al., 2023b), introduces a lightweight Q-Former designed to bridge modality gaps and leverages frozen LLMs. Leveraging LLMs, BLIP-2 can conduct zero-shot image-to-text generation using natural language prompts. We select the BLIP2-OPT-6.7B version for evaluation. [^2] https://huggingface.co/Salesforce/blip2-opt-6.7b
InstructBLIP (Dai et al., 2023) employs an instruction-aware visual feature extraction module based on BLIP2 (Li et al., 2023b) and is trained with unified multimodal instruction tuning datasets. We choose InstructBLIP-Vicuna-7B for evaluation. [^3] https://huggingface.co/Salesforce/instructblip-vicuna-7b.
LLaVA (Liu et al., 2023b), adopts Vicuna models as the backbone LLM and is trained on the ChatGPT/GPT-4 generated instruction tuning dataset. LLaVA-v1.5 (Liu et al., 2023a) improves on LLaVA models by employing curated task datasets and an enhanced modality alignment module. We evaluate both LLaVA-v1.5-7B [^4] https://huggingface.co/liuhaotian/llava-v1.5-7b and LLaVA-v1.5-13B. [^5] https://huggingface.co/liuhaotian/llava-v1.5-13b
Flamingo (Alayrac et al., 2022) pioneers the development of LVLMs by introducing a cross-gated layer for LLMs to produce visual-grounded text. The training dataset consists of interleaved visual data and text from the web pages, enabling it to generate free-form text as the output. We select the open-source implementation OpenFlamingo-9B (Awadalla et al., 2023) for evaluation. [^6] https://huggingface.co/openflamingo/OpenFlamingo-9B-vitl-mpt7b
IDEFICS is another open-sourced implementation of Flamingo (Alayrac et al., 2022) Trained on publicly available image-text alignment pairs and instruction tuning datasets, it demonstrates comparable results with the original closed-source model on various image-text benchmarks. We select the IDEFICS-Instruct-9B for evaluation. [^7] https://huggingface.co/HuggingFaceM4/idefics-9b-instruct.
Qwen-VL-Chat (Bai et al., 2023b) is a bilingual LVLM that supports both English and Chinese built on the Qwen LLM (Bai et al., 2023a). During the training phase, Qwen-VL-Chat adopts a packing strategy to create multiple images as inputs, improving its ability to understand the vision context. We select the Qwen-VL-Chat-7B for evaluation. [^8] https://github.com/QwenLM/Qwen-VL
GPT-4V  (OpenAI, 2023), the proprietary vision-language models developed by OpenAI, which are shown to be powerful on various multi-modal tasks (Yang et al., 2023a). The API version we queried is gpt-4-vision-preview.
Bard (Google, 2023) (now Gemini), a commercial LVLM developed by Google. We utilize the unofficial API [^9] https://github.com/dsdanielpark/Bard-API querying the model with our task prompts, accessed on 2023-11-17.
B.2 Task Prompts
We evaluate all the models with the same task prompts in our experiments, and the prompts for our four tasks are listed below:
  • Single-Figure Captioning: Create a caption for the provided figure.
  • Multiple-Figure Captioning: Create a caption for the provided figures.
  • Contextualized Captioning: We reuse the prompts in previous captioning tasks depending on the current figure type.
  • Title Generation: According to the figures and captions, generate a title for this paper. Title:
Appendix C Error Analysis
C.1 Caption Type Illustration
We illustrate captions of four types in our main paper in Figure 14. The Acceptable caption provides a comprehensive description of the figure presented. The oversimplified caption is too short compared with the original human-written caption. Furthermore, as show in the third block in Figure 14, Contextual Misinterpretation refers to captions with unmentioned content in the figure, such as the dataset colored in red. Recognition Error denotes the model wrongly identified the number or text in the figure, such as the misidentified model name in the last block of Figure 14.
C.2 Failure Sample of MathVista
Figure 15 shows a challenging geometry mathematic reasoning problem where both models fail to produce the correct answer. Echoing the quantitative results in our main paper, we believe future studies can incorporate more focused corpus for enhancing the geometry and mathematical reasoning ability of LVLMs.
Figure 14: Illustration of acceptable and three error types of generated captions.
Figure 15: A failure case on the geometry problem-solving task.